OpenStudy (anonymous):

Find the equation of the plane through (1,1,1) and containing the line which is the intersection of the planes x-y=2 and y-z=1

OpenStudy (anonymous):

This is what i think should be done..

OpenStudy (anonymous):

I got the vectors perpendicular to the gives planes, namely: v=i-j and u=j-k, then i took the cross product. This resulant vector is parrallel to the line of intersection. Now i believe i am in a position to write the equation of the plane since i have a vector perpinducalr to the plane and a point